# Nearly-zero large-angle anisotropy of the cosmic microwave background

### v5

@@ 3.2 Causal bounds on coherence and correlation
Inflation solves the main problem by introducing early cosmic acceleration, so that the comoving causal horizon moves closer with time rather than farther away  [^1] An excellent tutorial including visualizations of inflationary conformal space-time, causal horizons, and sky projections can be found at https://www.astro.ucla.edu/~wright/cosmo_04.htm. If the scale factor $a(t)$ undergoes many orders of magnitude of expansion during early acceleration with $\ddot{a}>0$ before some epoch $\eta=0$, even very distant comoving world lines were once in causal contact. This causal relationship is shown in Fig. (2): a comoving world line at any finite radial distance $\eta$ lies within the past light cone or “inflationary horizon” ${\cal H}$ at times earlier than $-\eta$. Fig. (2) shows spatial “footprints” of horizons: comoving spherical surfaces ${\cal S}(\eta)$ that pass through the horizon and out of causal contact at time $-\eta$, and come back into view after inflation at time $+\eta$. All points on a spherical surface ${\cal S}(\eta)$ have a causal connection with an event at its center at $-\eta$, so large-scale homogeneity and isotropy, as assumed in Eq. (9), can in principle be generated by a causal physical process.
@@
Now consider the physical process that generates spatial departures from uniformity caused by quantum fluctuations. In the analysis below, we consider a new causal symmetry that follows from a stronger causal constraint than standard inflationary theory. Suppose that all gravitational quantum fluctuations are causally coherent, in a sense well established from direct measurements of quantum entanglement (Zeilinger, 1999; Vilasini & Renner, 2024): physical correlations are not generated by systems in separate regions of space-time. This constraint requires that physical effects of coherent quantum fluctuation states are spatially compact. Specifically, we posit that quantum curvature fluctuations on any world line interval are entangled nonlocally with other locations only within the compact 4-dimensional region of space-time encompassed by its causal diamond. Put another way, physical correlations are bounded by two-way causal relationships.
We further posit that the inflationary horizon ${\cal H}$ imprints a sharp boundary on coherent quantum fluctuations that create curvature perturbations correlated with that world line: Quantum fluctuations create differences in classical potential from any world line in the future of its inflationary horizon ${\cal H}$. That is, the classical potential difference between world lines forms when they cross each others’ horizons.
@@
As discussed in the Appendix, this hypothesis about how quantum fluctuations convert into classical perturbations differs physically from the freezing of fluctuations in the standard QFT model, where coherent plane waves freeze independently on each comoving scale by synchronous cooling as their wavelengths stretch beyond the horizon scale. In that picture, the final observed classical correlations on ${\cal S}(\eta)$ are fixed by initial data laid down coherently in a region much larger than $\eta$, at a time much earlier than $-\eta$.
@@
This formulation is conformally invariant, so the coherent causal constraint applies to relational perturbations on all comoving length scales. As discussed in the Appendix, there may be observable effects of exotic high-order correlations in the 3D pattern of classical curvature perturbations on smaller scales than the current horizon.
@@
Figure 3: A projection along the time dimension of the relationships in Figure 3. The blue and green disks are the ellipses at $\eta=0$. This projection illustrates the geometrical derivation of the disentanglement angle $\Theta_{D}$ (Eq. 12). The angular radius of the intersection circle on ${\cal S}_{A}(\eta_{0}/2)$, $\pi/2-\Theta_{D}/2$, is the maximum angular separation for correlation with external perturbations in any direction, so the directional average in the correlation function $C(\Theta)$ (Eq. 14) vanishes at $\Theta=90^{\circ}\pm\Theta_{D}/2$ (Eq. 15). At the same time, perturbation differences on ${\cal S}_{B}(\eta_{0})$ are independent of causal restrictions, as in the standard picture, at angular separations $<<\Theta_{D}$. These angular causal relationships are conformally invariant, so they apply to any location, epoch or comoving length scale $\eta_{0}$.
@@ 3.3 Scale-invariant angular boundaries of causal correlation
Let us suppose that fluctuations generated in separate causal diamonds are separate quantum systems, and do not correlate with each other. We will now analyze how this 4D relationship maps onto angular correlation, which could lead to finite regions of angular separation in which no correlations occur from entanglement between causal diamonds. We refer to a region of angular separation with no causal entanglement as a “causal shadow.”
It is useful to classify angular boundaries according to geometrical criteria determined by different applications of causal-diamond coherence, which map onto different observational tests. These angular causal relationships are conformally invariant, so they apply on all comoving length scales.
3.3.1 3D causal shadow
@@
As discussed above, suppose that causal correlations are bounded by causal diamonds and inflationary horizons. Fig. (2) shows causal diamonds around two world lines $A$ and $B$, with separation $\eta_{0}$. The causal diamonds on these world lines that begin after $-\eta_{0}$ are disentangled from the other: their fluctuations constitute separate quantum systems, and generate uncorrelated perturbations.
Slices of comoving causal-diamond boundary surfaces in 3D are shown in Fig. (3). The sphere ${\cal S}_{A}(\eta_{0}/2)$ is shown because it represents the smallest horizon footprint around $A$ whose fluctuations are entangled with $B$. Smaller causal diamonds around $A$ intersect ${\cal S}_{B}(\eta_{0})$ at larger angles from $B$, but their perturbations are independent of those at $B$. Thus, the circular intersection of ${\cal S}_{A}(\eta_{0}/2)$ with ${\cal S}_{B}(\eta_{0})$ lies at the largest angular separation from the $B$ direction of locations $\vec{r}$ where perturbations $\Phi(\vec{r})-\Phi(\vec{r}_{A})$ are causally correlated with $\Phi(\vec{r}_{B})-\Phi(\vec{r}_{A})$.
As shown in Fig. (3), the angular radius on ${\cal S}_{B}(\eta_{0})$ of its intersection with ${\cal S}_{A}(\eta_{0}/2)$ is
$\Theta_{D}=2\arcsin(1/4)\simeq 28.96^{\circ},$ (12)
and the angular radius of the same intersection on ${\cal S}_{A}(\eta_{0}/2)$ is
$\pi/2-\Theta_{D}/2.$ (13)
@@
Thus, as viewed from $A$, correlations with perturbations at $B$ vanish at angles separated from the $B$ direction $\vec{\Omega}_{B}$ by more than $\pi/2-\Theta_{D}/2$. That is, values of $\Phi-\Phi_{A}$ at locations closer than $B$ with larger angular separation than Eq. (13) are independent of those at $B$.
@@
The correlation function viewed at $A$ is given by Eq. (6), an all-sky average correlation of $\Phi_{B}$ with $\bar{\Phi}_{B\Theta}$, the average value in a circle of radius $\Theta$ centered on $\vec{\Omega}_{B}$:
$C(\Theta)=\langle\Phi_{B}\bar{\Phi}_{B\Theta}\rangle_{\vec{\Omega}_{B}}.$ (14)
@@
Independence of $\Phi$ at angular separation exceeding $\pi/2-\Theta_{D}/2$ from every direction $\vec{\Omega}_{B}$ then leads to zero angular correlation of correlation over a range of angles symmetric around $\pi/2$:
$C_{\mathrm{3D}}(|\Theta-\pi/2|<\arcsin[1/4])=0.$ (15)
As discussed below, this symmetry applies separately to both the even- and odd- parity components of correlation, which are respectively symmetric and antisymmetric around $\Theta=\pi/2$, as well as their sum. It holds for any value of $\eta_{0}$, so it applies to correlations in the interior of any sphere with points on its boundary.
Importantly for direct null tests of CMB anisotropy, Eq. (14) describes cross-correlation of perturbations $\Phi_{B}$ at $|\vec{r}-\vec{r}_{A}|=\eta_{0}$ with $\bar{\Phi}_{B\Theta}$ at $|\vec{r}-\vec{r}_{A}|<\eta_{0}$. The angular shadow of zero correlation derived from intersection of the two spheres at disentanglement (Eq. 15) applies to foreground perturbations correlated with the last scattering surface, and therefore to gravitational anisotropy generated via the ISW effect. The null symmetry applies to observable CMB correlation independent of the location or epoch of observation: even-parity correlation always vanishes over a range of angles $75.52^{\circ}<\Theta<104.48^{\circ}$.
The effect of causal coherence becomes smaller on small angular scales. At angular separations from $A$ on ${\cal S}_{B}(\eta_{0})$ smaller than $\Theta_{D}$, where angular spectra are generally used for cosmological tests, differences from $\Phi_{A}$ on ${\cal S}_{B}(\eta_{0})$ are independent of $\Phi_{A}-\Phi_{B}$, so perturbations approximate the causally-unconstrained independent phases assumed in the standard inflationary picture. As a result, causal constraints have a negligible effect on angular correlations at small scales, or on the 3D power spectrum for scales small compared to any averaging volume.
3.3.2 Thin-sphere causal shadow
Now consider angular correlations of points constrained to lie at the same comoving distance from $A$. At the time $-\eta_{0}$ when the $AB$ relationship is freezing out— when $B$ passes through ${\cal H}_{A}$— locations on ${\cal S}_{A}(\eta_{0})$ with angular separation from $B$ greater than $\Theta=\pi/3$ lie outside of ${\cal H}_{B}$. These points lie outside the causal diamond surface ${\cal S}_{B}(\eta_{0})$, so their angular causal correlation vanishes:
$C_{\rm tscs}(\Theta>\pi/3)=0.$ (16)
This correlation shadow does not have the same robust application to a precise CMB test as the narrower 3D shadow (Eq. 15): it corresponds to zero angular correlation of potentials only on a thin sphere, so it does not include all gravitational anisotropy from interior perturbations. However, it could define a causal bound on correlations indirectly generated with the antipodal direction, as discussed below.
@@ 3.3.3 Correlation at larger angular separation
As shown above, all points on a horizon footprint ${\cal S}_{A}$ have a causal connection with their center $A$. The measured $C(\Theta)$ represents an all-sky average (Eq. 14), so it is modified at all angles by perturbations of $A$ relative to ${\cal S}_{A}$. Subtraction of unobservable monopole and dipole components could indirectly generate apparent angular correlations outside the symmetric interval of the 3D causal shadow (Eq. 15).
We have not derived a model-independent causal symmetry for angular separations outside the 3D causal shadow. Even so, we should explore constraints on correlations that include both odd and even parity components, and that also account for the unobserved dipole, to test whether data is consistent with zero total true correlation over a larger range of angular separations. Since we aim to avoid model-dependent assumptions, the simplest approach is to simply exclude large angular separations. For one test, we will exclude separations within the thin-sphere causal bound for correlations with the antipodal point (Eq. 16), $\Theta>2\pi/3$:
@@
We will also test the possibility of a wider causal shadow,
@@ 4.1 Dipole subtraction and parity separation
It is not possible to measure the true primordial pattern in the CMB, because the dipole components $a_{1m}$ have been removed from the maps to compensate for the local motion relative to the local cosmic rest frame, including our nonlinear orbits within the galaxy and the Local Group. These motions are not known to nearly enough precision to separate the primordial dipole (Peebles, 2022). Nevertheless, a small fraction of the subtracted dipole is part of the intrinsic large-angle primordial pattern on spherical causal diamond surfaces and contributes to correlation in the angular range of causal shadows. Thus, a null shadow symmetry can only become apparent when the intrinsic portion of the dipole is included. If it is a true symmetry of the cosmic horizon, then there must exist a dipole that can be added to the observed CMB temperature map that realizes the symmetry.
The total correlation (Eq. 4) is a sum of even and odd Legendre polynomials, which are respectively symmetric and antisymmetric about $\Theta=\pi/2$. To produce zero correlation over a range symmetric around $\Theta=\pi/2$, no combination of even functions can cancel any combination of odd ones, so if an angular correlation function vanishes over a range $[\pi/2-\Theta_{0},\pi/2+\Theta_{0}]$ symmetric about $\Theta=\pi/2$, the even contributions and the odd contributions to the angular correlation function must vanish independently over that range.
This property allows a direct, model- and dipole- independent test of causal symmetry, that uses only even-parity correlation. In a band of angles symmetric around $\Theta=\pi/2$ determined by the 3D causal shadow (Eq. 15), the sum of even terms must vanish on its own, independently of any dipole or model parameters. The causal shadow of even-parity correlation is thus
@@
Furthermore, if the true primordial angular correlation including $\ell=1$ vanishes over an arbitrary range $[\alpha,\beta]$, then the sum of the even and odd Legendre polynomials in Eq. (4), measured only with $\ell>1$, departs from zero by a function of known form, the dipole harmonic term
$\mathcal{D}(\Theta)=\frac{3}{4\pi}C_{1}\cos(\Theta),$ (20)
where $C_{1}\geq 0$. Thus, if there is a causal shadow over a larger angular range (Eq. 17 or 18), the sum of the even and odd Legendre polynomials for $\ell>1$ must vanish after addition of a dipole of unknown amplitude.
@@ 4.2.2 Standard realizations
Standard realized correlation functions include only $\ell>1$ harmonics. For the comparisons of total correlation shown in the second panel of Fig. (1), which include odd-parity harmonics, each realization is modified with a function of the form in Eq. (20) to minimize its residuals from zero. For realizations, this term does not have any relation to an intrinsic physical dipole: it is a “mock dipole” added to estimate how frequently the sum of $\ell>1$ harmonics in Eq. (4) comes as close to the maps as zero correlation in the posited range of angular separation. Although the residual variance of these comparisons exceeds that of purely even parity correlation in the 3D causal shadow, it appears that the measured $C(\Theta)$ in the posited range (Eqs. 17 or 18) is still closer to zero than almost all standard realizations, even when they have a mock dipole added.
@@ 4.2.3 Residuals
where $\ell_{\text{max}}=30$. Let $\{\Theta_{j,[\alpha,\beta]}\}_{i=1}^{N}$ denote a uniformly spaced lattice of points in the range $[\alpha,\beta]$. Then, define the even-parity residual
[…]
$\approx\sum_{i=1}^{N}[C_{\text{\rm even}}(\Theta_{i,[\alpha,\beta]})]^{2}\cdot\left(\frac{\beta-\alpha}{N}\right)\!,.$ (23)
@@
[…]
$\approx\sum_{i=1}^{N}[\tilde{C}_{\beta}(\Theta_{i,[\alpha,\beta]})]^{2}\cdot\left(\frac{\beta-\alpha}{N}\right)\!,$ (25)
@@
$\Delta_{\text{\rm even}}$ $\equiv\Delta_{\text{\rm even},[\pi/2-\Theta_{D}/2,\pi/2+\Theta_{D}/2]},$ (26)
$\Delta_{\text{minimal}}$ $\equiv\Delta_{\text{best-fit},[\pi/2-\Theta_{D}/2,2\pi/3]},$ (27)
$\Delta_{\text{maximal}}$ $\equiv\Delta_{\text{best-fit},[\pi/2-\Theta_{D}/2,3\pi/4]},$ (28)
where $\Theta_{D}/2=\arcsin(1/4)$ as derived above (Eq. 12). We use these three residuals as a measure of how compatible a given power spectrum $\left\{C_{\ell}\right\}_{\ell>1}$ is with the causal shadow symmetry. Each of the integrals must vanish for a power spectrum that exactly agrees with the causal shadow symmetry in the specified range. In practice, we found that $N=2000$ is a sufficiently high lattice resolution to approximate the integrals among different data sets and standard model realizations with negligible error.
Figure 4: Cumulative probability of deviations from zero correlation in standard realizations, compared with deviations of CMB maps, over various ranges of angular separation. The top panel directly compares deviations of even-parity correlation (Eq. 23) over the computed 3D causal shadow (Eq. 19). No parameters are used for this comparison. The middle panel compares deviations of total correlation (Eq. 25) over the minimal asymmetric shadow (Eq. 17), and the bottom panel compares deviations of total correlation over the maximal asymmetric shadow (Eq. 18); both of these allow for a mock-dipole correction to minimize residuals for each realization. In spite of variation between the maps, their variations are all much closer to zero than almost all standard realizations. The departures from zero are comparable with the differences between the maps, as expected if they are dominated by systematic measurement errors. In all of these comparisons, the Planck maps match zero better than the WMAP maps. In the cleanest direct comparison, which is represented by the top panel, the residuals $\Delta_{\text{\rm even}}$ found in the Planck maps are only 1 to 9 $\mu{\rm K}^{4}$, compared with $\Delta_{\text{\rm even}}\simeq 10^{4}\mu{\rm K}^{4}$ found in typical realizations. The probabilities for the standard picture to match such small values range from $10^{-4.3}$ to $10^{-2.8}$. The extraordinary overall precision of CMB isotropy can be characterized by the fractional measured dimensionless residual, which reaches its smallest value $\int C^{2}_{\rm even}/(2.7{\rm K})^{4}\sim 2\times 10^{-26}$ in the NILC map.

### v7

@@ 4.2 Causal bounds on coherence and correlation
[…]
@@
Now consider the physical process that generates spatial departures from uniformity caused by quantum fluctuations. In the analysis below, we consider a new causal symmetry that follows from a stronger causal constraint than standard inflationary theory. Suppose that all gravitational quantum fluctuations are causally coherent, in a sense well established from direct measurements of quantum entanglement (Zeilinger, 1999; Vilasini & Renner, 2024): physical correlations are not generated by systems in separate regions of space-time. This constraint requires that physical effects of coherent quantum fluctuation states are spatially compact. Specifically, we posit that quantum curvature fluctuations on any world line interval are entangled nonlocally with other locations only within the compact 4-dimensional region of space-time encompassed by its causal diamond. Put another way, physical correlations are bounded by two-way causal relationships.  [^5] Effects of causally coherent fluctuations with the same physical origin in flat space-time may be detectable in proposed laboratory experiments (Vermeulen et al., 2021; Kwon, 2025; Vermeulen et al., 2025).
[…]
@@
As discussed in the Appendix, this hypothesis about how quantum fluctuations convert into classical perturbations differs physically from the freezing of fluctuations in the standard QFT model, where coherent plane waves freeze independently on each comoving scale by synchronous cooling as their wavelengths stretch beyond the horizon scale. In that picture, the final observed classical correlations on ${\cal S}(\eta)$ are fixed by initial data laid down coherently in a region much larger than $\eta$, at a time much earlier than $\eta$. Perturbations in the standard picture are defined in relation to a fixed initial background; in a causally coherent model, they are defined relationally between world lines, as allowed by causal relationships in a coherent quantum system. In standard inflation, the final outcome everywhere is fixed by state of the initial vacuum at the start of inflation; in a causally coherent picture, the quantum state and the final relational perturbations remain in a superposition within horizons until the end of inflation.
@@
Our formulation is conformally invariant, so the coherent causal constraint applies to relational perturbations on all comoving length scales. As discussed in the Appendix, there may be observable effects of exotic high-order correlations in the 3D pattern of classical curvature perturbations on smaller scales than the current horizon.
@@
Figure 3: A section of relationships in Figure 2 at time $\eta=0$. The blue and green disks represent equatorial slices of comoving spheres at $\eta=0$, shown as ellipses in Figure 2. This projection illustrates the geometrical derivation of the disentanglement angle (Eq. 12). The angular radius of the intersection circle on ${\cal S}_{A}(\eta_{0})$, $\theta_{d}=\pi/2-\arcsin(1/4)$, is the maximum angular separation for correlation with any direction, so the directional average in the correlation function $C(\Theta)$ (Eq. 13) vanishes at $\Theta=90^{\circ}\pm\arcsin(1/4)$ (Eq. 15). At the same time, perturbation differences from $\Phi_{B}$ on ${\cal S}_{B}(2\eta_{0})$ are independent of causal restrictions, as in the standard picture, at angular separations $\lesssim 2\arcsin(1/4)\simeq 29^{\circ}$. These angular causal relationships are conformally invariant, so they apply to any location, epoch or comoving length scale $\eta_{0}$.
@@ 4.3 Scale-invariant angular boundaries of causal correlation
4.3.1 Zero correlation around $\Theta=90^{\circ}$ from disentanglement
@@
As discussed above, suppose that causal correlations are bounded by causal diamonds and perturbations form on inflationary horizons. Fig. (2) shows causal diamonds around two world lines $A$ and $B$, with separation $2\eta_{0}$. The causal diamonds on these world lines that begin after $2\eta_{0}$ are disentangled from the other: their fluctuations constitute separate quantum systems, and generate uncorrelated perturbations on their respective horizons.
These world lines are shown because the circular intersection of ${\cal S}_{A}(\eta_{0})$ with ${\cal S}_{B}(2\eta_{0})$ is the largest angle from the $AB$ axis where $\Phi$ at $|\vec{r}|\leq\eta_{0}$ is correlated by entanglement with horizons centered at any location on the axis. (Horizons of points $B$ at larger distances from $A$ intersect at larger angles, but they are not entangled within ${\cal S}_{A}(\eta_{0})$.)
Slices of the corresponding comoving causal-diamond boundary surfaces in 3D are shown in Fig. (3). The angular radius on ${\cal S}_{A}(\eta_{0})$ of its intersection with ${\cal S}_{B}(2\eta_{0})$ is the disentanglement angle,
$\theta_{d}=\pi/2-\arcsin(1/4)\simeq 75.52^{\circ}.$ (12)
@@
As viewed from $A$ at time $\eta_{0}$, perturbations separated from the $B$ direction $\vec{\Omega}_{B}$ by more than $\theta_{d}$ are independent of those at $B$.
@@
The angular correlation function is given by Eq. (6), an all-sky correlation with circles of radius $\Theta$ centered on every direction $\vec{\Omega}_{B}$:
$C(\Theta)=\langle\Phi_{\vec{\Omega}_{B}}\bar{\Phi}_{\vec{\Omega}_{B}\Theta}\rangle_{\vec{\Omega}_{B}}.$ (13)
@@
Independence of $\Phi$ at polar separation exceeding the disentanglement angle (Eq. 12) from every direction $\vec{\Omega}_{B}$ then leads to zero correlation above the disentanglement angle,
$C(\Theta>\theta_{d})=0.$ (14)
Importantly for direct null tests of CMB anisotropy, the bound includes angular cross correlations at different radii (of $\Phi_{\vec{\Omega}_{B}}(|\vec{r}|<\eta_{0})$ with $\bar{\Phi}_{\vec{\Omega}_{B}\Theta}(|\vec{r}|=\eta_{0})$), so it applies to gravitational anisotropy generated via the ISW effect as well as from last scattering.
The even- and odd- parity components of $C(\Theta$ are respectively symmetric and antisymmetric around $\Theta=\pi/2$, so they must both have zero correlation over a range of angles symmetric around $\pi/2$:
$C_{\mathrm{total}}=C_{\mathrm{even}}=C_{\mathrm{odd}}=0\ \qquad(75.52^{\circ}<\Theta<104.48^{\circ}).$ (15)
As described below, because odd parity correlations always vanish at exactly $90^{\circ}$ and the dipole component is not measured, the even-parity component provides the most powerful direct test of this “causal shadow” symmetry.
@@ 4.3.2 Correlation at larger angular separation
At angles larger than $\pi/2+\arcsin(1/4)\simeq 104.48^{\circ}$, tests of null symmetry (Eq. 14) must include odd-parity correlation. Moreover, where odd and even parity components do not both vanish, subtraction of unmeasured monopole and dipole components can indirectly generate apparent angular correlations outside the symmetric interval of the 3D causal shadow (Eq. 15).  [^6] Apparent large-angle correlation introduced by monopole and dipole subtraction was studied in the context of wavelet models by Copi et al. (2019).
We have not derived a model-independent causal symmetry for angular separations outside the 3D causal shadow. Even so, it is interesting to explore empirical constraints on correlations that include both odd and even parity components, and that also account for the unobserved dipole, to test whether data is consistent with zero total true correlation over a larger range of angular separations. We choose to test two ranges with possible geometrical origins. For one test, we will exclude separations within the intersection angle of a horizon of equal radius centered on the antipodal point, $\Theta>2\pi/3$:
@@
We will also test the possibility of a wider causal shadow that could be generated by causally-coherent “tilted” perturbations:
@@ 5.1 Dipole subtraction and parity separation
It is not possible to measure the true primordial pattern in the CMB, because the dipole components $a_{1m}$ have been removed from the maps to compensate for the local motion relative to the local cosmic rest frame, including our nonlinear orbits within the galaxy and the Local Group. These motions are not known to nearly enough precision to separate the primordial dipole (Peebles, 2022). Nevertheless, a small fraction of the subtracted dipole is part of the intrinsic large-angle primordial pattern on spherical causal diamond surfaces and contributes to correlation in the angular range of causal shadows. Thus, a null shadow symmetry can only become apparent when the intrinsic portion of the dipole is included. If it is a true symmetry of gravitational potential correlation, then there must exist a dipole that can be added to the observed CMB temperature map that realizes the symmetry.  [^7] We omit consideration of second-order Doppler anisotropy, which generates even-parity harmonics including a quadrupole, but with amplitude smaller than our measurement precision.
The total correlation (Eq. 4) is a sum of even and odd Legendre polynomials, which are respectively symmetric and antisymmetric about $\Theta=\pi/2$. To produce zero correlation over a range symmetric around $\Theta=\pi/2$, no combination of even functions can cancel any combination of odd ones, so if an angular correlation function vanishes over a range $\pi/2-\Theta_{0},\pi/2+\Theta_{0}]$ symmetric about $\Theta=\pi/2$, the even contributions and the odd contributions to the angular correlation function must vanish independently over that range.
This property allows a direct, model- and dipole- independent test of causal symmetry, that uses only even-parity correlation. In a band of angles symmetric around $\Theta=\pi/2$ determined by the causal shadow (Eq. 15), the sum of even terms must vanish on its own, independently of any dipole or model parameters. The causal shadow of even-parity correlation is thus
@@
Furthermore, if the true primordial angular correlation including $\ell=1$ vanishes over an arbitrary range $\alpha,\beta]$, then the sum of the even and odd Legendre polynomials in Eq. (4), measured only with $\ell>1$, departs from zero by a function of known form, the dipole harmonic term
$\mathcal{D}(\Theta)=C_{\rm dipole\ only}(\Theta)=\frac{3}{4\pi}C_{1}\cos(\Theta),$ (19)
where $C_{1}\geq 0$. Thus, if there is a causal shadow over a larger angular range (Eq. 16 or 17), the sum of the even and odd Legendre polynomials for $\ell>1$ must vanish after addition of a dipole of unknown amplitude.
@@ 5.2.2 Standard realizations
Standard realized correlation functions include only $\ell>1$ harmonics. For the comparisons of total correlation shown in the second panel of Fig. (1), which include odd-parity harmonics, each realization is modified with a function of the form in Eq. (19) to minimize its residuals from zero. For realizations, this term does not have any relation to an actual physical dipole: it is a “mock dipole” added to estimate how frequently the sum of $\ell>1$ harmonics in Eq. (4) comes as close to the maps as zero correlation in the posited range of angular separation, even if a dipole of unrestricted value is included. [^8] Our likelihood estimates are conservatively generous to the standard picture; we have not allowed for the fact that most realizations do not have dipoles as large as the mock dipoles. Although the residual variance of these comparisons exceeds that of purely even parity correlation in the 3D causal shadow, it appears that the measured $C(\Theta)$ in the posited range (Eqs. 16 or 17) is still closer to zero than almost all standard realizations, even when they have a mock dipole added.
@@ 5.2.3 Residuals
where $\ell_{\text{max}}=30$. Let $\{\Theta_{j,[\alpha,\beta]}\}_{i=1}^{N}$ denote a uniformly spaced lattice of points in the range $\alpha,\beta]$. Then, define the even-parity residual
$\Delta_{\text{\rm even},[\alpha,\beta]}(C(\Theta))$ $\equiv\int_{\alpha}^{\beta}|C_{\text{\rm even}}(\Theta)|^{2}\,d\Theta\>\!$ (21)
$\hskip-30.0pt\approx\sum_{i=1}^{N}[C_{\text{\rm even}}(\Theta_{i,[\alpha,\beta]})]^{2}\cdot\left(\frac{\beta-\alpha}{N}\right)\!.$ (22)
@@
$\Delta_{\text{best-fit},[\alpha,\beta]}(C(\Theta))$ $\equiv\int_{\alpha}^{\beta}|\tilde{C}_{\beta}(\Theta)|^{2}\,d\Theta\>\!$ (23)
$\hskip-30.0pt\approx\sum_{i=1}^{N}[\tilde{C}_{\beta}(\Theta_{i,[\alpha,\beta]})]^{2}\cdot\left(\frac{\beta-\alpha}{N}\right)\!,$ (24)
@@
$\Delta_{\text{\rm even}}$ $\equiv\Delta_{\text{\rm even},[\pi/2-\arcsin(1/4),\ \pi/2+\arcsin(1/4)]},$ (25)
$\Delta_{\text{minimal}}$ $\equiv\Delta_{\text{best-fit},[\pi/2-\arcsin(1/4),\ 2\pi/3]},$ (26)
$\Delta_{\text{maximal}}$ $\equiv\Delta_{\text{best-fit},[\pi/2-\arcsin(1/4),\ 3\pi/4]}.$ (27)
using the angular relationships discussed above (Fig. 3). We use these three residuals as a measure of how compatible a given power spectrum $\left\{C_{\ell}\right\}_{\ell>1}$ is with the causal shadow symmetry. Each of the integrals must vanish for a power spectrum that exactly agrees with the causal shadow symmetry in the specified range. In practice, we found that $N=2000$ is a sufficiently high lattice resolution to approximate the integrals among different data sets and standard model realizations with negligible error.
Figure 4: Cumulative probability of deviations from zero correlation in standard realizations, compared with deviations of CMB maps, over various ranges of angular separation. The top panel directly compares deviations of even-parity correlation (Eq. 22) over the computed 3D causal shadow (Eq. 18). No parameters are used for this comparison. The middle panel compares deviations of total correlation (Eq. 24) over the minimal asymmetric shadow (Eq. 16), and the bottom panel compares deviations of total correlation over the maximal asymmetric shadow (Eq. 17); both of these allow for a mock-dipole correction to minimize residuals for each realization. In spite of variation between the maps, their variations are all much closer to zero than almost all standard realizations. The departures from zero are comparable with the differences between the maps, as expected if they are dominated by systematic measurement errors. In all of these comparisons, the Planck maps match zero better than the WMAP maps. In the cleanest direct comparison, which is represented by the top panel, the residuals $\Delta_{\text{\rm even}}$ found in the Planck maps are only 1 to 9 $\mu{\rm K}^{4}$, compared with $\Delta_{\text{\rm even}}\simeq 10^{4}\mu{\rm K}^{4}$ found in typical realizations. The probabilities for the standard picture to match such small values range from $10^{-4.3}$ to $10^{-2.8}$.